# Syntomic cohomology of truncated Brown–Peterson spectra

### v1

@@ Abstract.
We compute the $\operatorname{MU}$-based syntomic cohomologies, mod $(p,v_{1},\cdots,v_{n})$, of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of the truncated Brown–Peterson spectrum $\operatorname{BP}\langle n\rangle$. As qualitative consequences, we resolve the Lichtenbaum–Quillen, telescope, and redshift questions for the algebraic K-theories of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$. This extends work of the Hahn and Wilson. We also explicitly compute the algebraic K-theory of arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$ at all primes $p\geq 5$ extending previous work of the author, Ausoni, Culver, Höning, and Rognes.
@@ 1. Introduction
By celebrated work of Quillen [QUI69], complex cobordism provides a bridge between geometric topology and algebraic geometry. Working $p$-locally, a homology theory called Brown–Peterson homology $\mathrm{BP}$ splits off of complex cobordism. This homology theory has coefficients a polynomial algebra over the $p$-local integers with generators $v_{i}$ for $i\geq 1$. Complex oriented homology theories carry a formal group law, and the complex oriented Brown–Peterson homology theory carries the universal $p$-typical formal group law. By Lazard [LAZ55], $p$-typical (1-dimensional) formal group laws are classified by their height and we may ask whether there is a complex oriented homology theory that carries a formal group law of height $\leq n$. One way to build such a ring spectrum is to take a quasi-regular quotient of Brown–Peterson homology to produce the truncated Brown-Peterson homology $\mathrm{BP}\langle n\rangle$ with coefficients a polynomial algebra over the $p$-local integers on generators $v_{1}$, $v_{2}$, $\dots$ $v_{n}$.
@@
Structured ring spectra play a fundamental role in homotopy theory, for example more highly structured ring spectra are equipped with a richer notion of power operations. It is known that complex cobordism is a $\operatorname{\mathbb{E}}_{\infty}$-ring, which means it is a homology theory equipped with a coherent commutative multiplication. It was hoped that the Brown–Peterson homology theory and the truncated Brown–Peterson cohomology theories would have similar structure. It was shown that $\operatorname{\mathbb{E}}_{\infty}$ algebra structures exists on $\operatorname{BP}\langle n\rangle$ for n=1,0,1n=-1,0,1 at all primes and $n=2$ at the primes $p=2,3$ by [BR08, HL10, LN12, LN14] and an $\operatorname{\mathbb{E}}_{4}$-algebra structure exists on $\operatorname{BP}$ by [BM13]. However, it was proven by [LAW18, SEN24] that $\operatorname{BP}$ and $\operatorname{BP}\langle n\rangle$ do not have $\operatorname{\mathbb{E}}_{\infty}$ algebra structures when $n\geq 4$, so we do not live in the best of all possible worlds.
@@
Producing $\operatorname{\mathbb{E}}_{m}$ algebra structures on truncated Brown–Peterson spectra for $m>1$ at arbitrary heights $n$ remained an open problem until [HW22] proved that there exist $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra for all primes $p$ and heights $n$. These forms depend on a choice of generators $v_{i}$. The most common choice would be to choose the Araki or Hazewinkle generators. However, it was observed by Strickland [STR99, Remark 6.5] at the prime $2$ that, although these forms of truncated Brown–Peterson spectra are equipped with $\operatorname{\mathbb{E}}_{1}$ algebra structures, they are not equipped with additional structure. In particular, they are not equipped with $\operatorname{\mathbb{E}}_{2}$ algebra structures.
@@
To compute invariants of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra we are therefore forced to work without additional structure. The algebraic K-theory and topological Hochschild homology of an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra is only an $\operatorname{\mathbb{E}}_{0}$ algebra over the corresponding invariant for $\operatorname{MU}$, where an $\operatorname{\mathbb{E}}_{0}$ algebra simply consists of a unit map. Nevertheless, the fact that the unit map $\operatorname{MU}_{(p)}\to\operatorname{BP}\langle n\rangle$ is a $\pi_{*}$-surjection allows us to proceed. We therefore resolve the redshift, Lichtenbaum–Quillen and telescope questions in the case of algebraic K-theory of truncated Brown–Peterson spectra. Let $T(n)=\mathbb{S}/(p^{i_{0}},\cdots,v_{n-1}^{i_{n-1}})[v_{n}^{-1}]$ for suitably large integers $i_{0},\cdots,i_{n-1}$. We write $L_{n}$ for the Bousfield localization at $\mathbb{Q}\oplus K(1)\oplus\cdots K(n)$ where $K(i)$ is the $i$-th Morava K-theory at the prime $p$. We also write $L_{n}^{f}$ for the Bousfield localization at $\mathbb{Q}\oplus T(1)\oplus\cdots\oplus T(n)$.
@@
Let $p$ be a prime number and $n\geq 1$ an integer. Let $\operatorname{BP}\langle n\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the $n$-th truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1. Then
@@
    (Lichtenbaum–Quillen) The localization map $\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$ has bounded above fiber.
@@
Let $n\geq 1$ be an integer. The mod $(p,v_{1},\cdots,v_{n+1})$ syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
$\operatorname{\mathbb{F}}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle\oplus\bigoplus_{i=1}^{n+1}\operatorname{\mathbb{F}}_{p}\langle\lambda_{s}:s\in\langle n\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
as a bigraded $\operatorname{\mathbb{F}}_{p}$-vector space and the $v_{n+1}$-Bockstein spectral sequence collapses. Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ and $\langle n\rangle=\{1,2,\cdots,n\}$.
@@
Let $p\geq 5$. We can identify
@@
When $n=0,1$, Theorem B was proven by [LW22] and [HRW22] respectively. As a consequence, Hahn–Raksit–Wilson extend Theorem A and the $n=1$ analogue of Theorem C to the prime $p=3$ in [HRW22].
@@
$\pi_{*}\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow\pi_{*}L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$
@@ 1.1. Acknowledgements
The author is indebted to Jeremy Hahn and Dylan Wilson who contributed significantly to this project, but declined to be co-authors. Additionally, the author benefited from input from Christian Ausoni, Robert Burklund, Ishan Levy, Piotr Pstrągowski, Arpon Raksit, John Rognes, Andrew Senger, and Tristan Yang, either from direct communication or indirect communication though Jeremy Hahn and Dylan Wilson.
@@ 2. Hochschild homology
First, we define forms of $\operatorname{BP}\langle n\rangle$ in Section 2.1 and the $\operatorname{MU}$-based motivic filtration in Section 2.2. We compute Hochschild homology of $\operatorname{MU}$, $\operatorname{BP}$, and $\operatorname{BP}\langle n\rangle$ with $\operatorname{\mathbb{F}}_{p}$-coefficients in Section 2.3 along with their motivic filtrations in Section 2.4.
@@ 2.1. Forms of $\operatorname{BP}\langle n\rangle$
By [CM15, Corollary 1.3], the Quillen idempotent defines an $\operatorname{\mathbb{E}}_{2}$ algebra map
@@
While $\pi_{*}\operatorname{BP}$ is only non-canonically isomorphic to a polynomial ring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots]$, for each $i\geq 1$ the subring $\mathbb{Z}_{(p)}[v_{1},\cdots,v_{i}]\subset\pi_{*}\operatorname{BP}$ is well-defined. Indeed, it is the subring generated by all elements of degree at most $2p^{i}-2$. Based on this, the following definition is standard:
@@
$\mathbb{Z}_{(p)}[v_{1},\cdots,v_{n}]\subset\operatorname{BP}_{*}\subset\pi_{*}\operatorname{MU}_{(p)}\to\pi_{*}R$
@@
In particular, the Eilenberg–MacLane spectrum $\mathbb{Z}_{(p)}$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 0\rangle$ and the Adams summand $\ell$ of $p$-local topological K-theory $\mathrm{ku}$ is a $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 1\rangle$.
@@ 2.2. The motivic filtration
We sometimes write $M/(p,\dots,v_{i})$ for the iterated tensor product
@@
$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(R)/(p,\cdots,v_{i}).$
@@
$M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}}(\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,\cdots,v_{i}))\cong M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}}(\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,\cdots,v_{i})).$
@@ 2.3. Hochschild homology with $\operatorname{\mathbb{F}}_{p}$-coefficients
of graded $\operatorname{\mathbb{F}}_{p}$-algebras, where $|\lambda_{i}^{\prime}|=2i+1$ by [MS93, Remark 4.3]. We write $\lambda_{j}\coloneqq\lambda_{p^{j}-1}^{\prime}$. The $\operatorname{\mathbb{E}}_{3}$ algebra $\operatorname{THH}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})$ is a retract of $\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ as an $\mathbb{E}_{1}$ algebra by [CM15, Corollary 1.3] and has homotopy groups
@@ 2.4. The motivic filtration on Hochschild homology
First, we need a preliminary result, which uses the notion of even flatness due to Pstrągowski, see [PST23, Definition 4.4].
Property 2.12.
The $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-module $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is an even flat. In particular, the spectrum $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$ is even.
@@
of $\operatorname{\mathbb{E}}_{0}$-$\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebras where
@@
By Proposition 2.12, we know that $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even. Therefore, the $p,v_{1},\cdots,v_{n}$-Bockstein spectral sequences all collapse and we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU})_{p}^{\wedge}$ is a finitely generated free $\operatorname{\mathbb{Z}}_{p}^{\wedge}[v_{1},\cdots v_{n}]$-module. Letting $B_{i}=\operatorname{BP}\langle i\rangle$ for $1\leq i\leq n$. We have cofiber sequences
@@
as well. Taking a further quotient by $v_{n+1}$ produces the result. To see this, consider the commutative diagram of $\operatorname{THH}(\operatorname{MU})$-modules
$\operatorname{MU}$$\operatorname{THH}(\operatorname{MU})$$\operatorname{BP}\langle n\rangle$$\operatorname{THH}(\operatorname{BP}\langle n\rangle)$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$}
to determine that the $v_{n+1}\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}$ acts trivially on $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ compatibly with the map $\operatorname{THH}(\operatorname{BP}\langle n\rangle)\to\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$. ∎

### v2

@@ Abstract.
We compute the $\operatorname{MU}$-based syntomic cohomologies, mod $(p,v_{1},\cdots,v_{n})$, of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of the truncated Brown–Peterson spectrum $\operatorname{BP}\langle n\rangle$. As qualitative consequences, we resolve the Lichtenbaum–Quillen, telescope, and redshift questions for the algebraic K-theories of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$. This extends work of Hahn and Wilson. We also explicitly compute the mod $(p,v_{1},v_{2})$ algebraic K-theory of arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$ at all primes $p\geq 5$ extending previous work of the author, Ausoni, Culver, Höning, and Rognes.
@@ 1. Introduction
By celebrated work of Quillen [QUI69], complex cobordism provides a bridge between geometric topology and algebraic geometry. Working $p$-locally, a homology theory called Brown–Peterson homology $\mathrm{BP}$ splits off of complex cobordism. This homology theory has coefficients a polynomial algebra over the $p$-local integers with generators $v_{i}$ for $i\geq 1$. Complex oriented homology theories carry a formal group law, and the complex oriented Brown–Peterson homology theory carries the universal $p$-typical formal group law. By Lazard [LAZ55], $p$-typical (1-dimensional) formal group laws are classified by their height and we may ask whether there is a complex oriented homology theory that carries a formal group law of height $\leq n$. One way to build such a ring spectrum is to take a quasi-regular quotient of Brown–Peterson homology to produce truncated Brown-Peterson homology $\mathrm{BP}\langle n\rangle$ with coefficients a polynomial algebra over the $p$-local integers on generators $v_{1}$, $v_{2}$, $\dots$ , $v_{n}$. By convention $\mathrm{BP}\langle-1\rangle=\mathbb{F}_{p}$.
@@
Structured ring spectra play a fundamental role in homotopy theory, for example more highly structured ring spectra are equipped with a richer notion of power operations. It is known that complex cobordism is a $\operatorname{\mathbb{E}}_{\infty}$-ring, which means it is a homology theory equipped with a coherent commutative multiplication. It was hoped that the Brown–Peterson homology theory and the truncated Brown–Peterson cohomology theories would have similar structure. It was shown that $\operatorname{\mathbb{E}}_{\infty}$-ring structures exist on $\operatorname{BP}\langle n\rangle$ for n=1,0,1n=-1,0,1 at all primes and $n=2$ at the primes $p=2,3$ by [BR08, HL10, LN12, LN14] and an $\operatorname{\mathbb{E}}_{4}$-ring structure exists on $\operatorname{BP}$ by [BM13]. However, it was proven by Lawson and Senger [LAW18, SEN24] that $\operatorname{BP}$ and $\operatorname{BP}\langle n\rangle$ do not have $\operatorname{\mathbb{E}}_{\infty}$-ring structures when $n\geq 4$, so we do not live in the best of all possible worlds.
@@
Producing $\operatorname{\mathbb{E}}_{m}$-ring structures on truncated Brown–Peterson spectra for $m>1$ at arbitrary heights $n$ remained an open problem until Hahn and Wilson [HW22] proved that there exist $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra for all primes $p$ and heights $n$. These forms depend on a choice of generators $v_{i}$. The most common choice would be to choose the Araki or Hazewinkle generators. However, it was observed by Strickland [STR99, Remark 6.5] at the prime $2$ that, although these forms of truncated Brown–Peterson spectra are equipped with $\operatorname{\mathbb{E}}_{1}$-ring structures, they are not equipped with additional structure. In particular, they are not equipped with $\operatorname{\mathbb{E}}_{2}$-ring structures.
@@
To compute invariants of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra we are therefore forced to work without additional structure. The algebraic K-theory and topological Hochschild homology of an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra is only an $\operatorname{\mathbb{E}}_{0}$ algebra over the corresponding invariant for $\operatorname{MU}$, where an $\operatorname{\mathbb{E}}_{0}$ algebra simply consists of a unit map. Nevertheless, the fact that the unit map $\operatorname{MU}_{(p)}\to\operatorname{BP}\langle n\rangle$ is a $\pi_{*}$-surjection allows us to proceed. We therefore resolve the redshift, Lichtenbaum–Quillen and telescope questions in the case of algebraic K-theory of truncated Brown–Peterson spectra. Let $T(n)=\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}}\cdots,v_{n-1}^{i_{n-1}})[v_{n}^{-1}]$ for suitably large integers $i_{0},i_{1},\cdots,i_{n-1}$. We write $L_{n}$ for the Bousfield localization at $\mathbb{Q}\oplus K(1)\oplus\cdots K(n)$ where $K(i)$ is the $i$-th Morava K-theory at the prime $p$. We also write $L_{n}^{f}$ for the Bousfield localization at $\mathbb{Q}\oplus T(1)\oplus\cdots\oplus T(n)$.
@@
Let $p$ be a prime number and $n\geq-1$ be an integer. Let $\operatorname{BP}\langle n\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the $n$-th truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1. Then
@@
    (Lichtenbaum–Quillen) The localization map $\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$ has bounded above fiber.
@@
Let $n\geq-1$ be an integer. The mod $(p,v_{1},\cdots,v_{n+1})$ syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
$\operatorname{\mathbb{F}}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle\oplus\bigoplus_{i=1}^{n+1}\operatorname{\mathbb{F}}_{p}\langle\lambda_{s}:s\in\langle n+1\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
as a bigraded $\operatorname{\mathbb{F}}_{p}$-vector space and the $v_{n+1}$-Bockstein spectral sequence collapses. Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ and $\langle n+1\rangle=\{1,2,\cdots,n+1\}$.
@@
Let $p\geq 5$ and let $\operatorname{BP}\langle 2\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the second truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1. We can identify
@@
When $n=0,1$, Theorem B was proven by [LW22] and [HRW22] respectively. As a consequence, Hahn–Raksit–Wilson [HRW22] extend Theorem A and the $n=1$ analogue of Theorem C to the prime $p=3$.
@@
$\pi_{*}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow\pi_{*}L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$
@@ 1.1. Acknowledgements
The author is indebted to Jeremy Hahn and Dylan Wilson who contributed significantly to this project, but declined to be co-authors. Additionally, the author benefited from input from Christian Ausoni, Robert Burklund, Ishan Levy, Piotr Pstrągowski, Arpon Raksit, John Rognes, Andrew Senger, and Tristan Yang, either from direct communication or indirect communication though Jeremy Hahn and Dylan Wilson. The author would also like to thank Shay Ben-Moshe for helpful feedback on a previous draft. The author is also grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support.
@@ 2. Hochschild homology
First, we define forms of $\operatorname{BP}\langle n\rangle$ in Section 2.1 and the $\operatorname{MU}$-based motivic filtration in Section 2.2. We compute Hochschild homology of $\operatorname{BP}\langle n\rangle$ with $\operatorname{\mathbb{F}}_{p}$-coefficients in Section 2.3 along with its motivic filtrations in Section 2.4.
@@ 2.1. Forms of $\operatorname{BP}\langle n\rangle$
By [CM15, Corollary 1.3], the Quillen idempotent defines an $\operatorname{\mathbb{E}}_{2}$-ring map
@@
While $\pi_{*}\operatorname{BP}$ is only non-canonically isomorphic to a polynomial ring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots]$, for each $i\geq 1$ the subring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots,v_{i}]\subset\pi_{*}\operatorname{BP}$ is well-defined. Indeed, it is the subring generated by all elements of degree at most $2p^{i}-2$. Based on this, the following definition is standard:
@@
$\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots,v_{n}]\subset\operatorname{BP}_{*}\subset\pi_{*}\operatorname{MU}_{(p)}\to\pi_{*}R$
@@
In particular, the Eilenberg–MacLane spectrum $\mathbb{Z}_{(p)}$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 0\rangle$ and the Adams summand $\ell$ of $p$-local topological K-theory $\mathrm{ku}$ is a $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 1\rangle$. By convention $\mathbb{F}_{p}$ is a $\mathbb{E}_{\infty}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle-1\rangle$.
@@
Convention 2.3.
Throughout the paper, whenever a prime $p$, the notations $L_{n}$, $L_{n}^{f}$ and $\operatorname{BP}\langle n\rangle$ appear in a statement then the implicit primes are always the same and agree with the given prime $p$ if it appears.
@@ 2.2. The motivic filtration
We sometimes write $M/(p,v_{1},\cdots,v_{i})$ for the iterated tensor product
@@
$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(R)/(p,v_{1},\cdots,v_{i}).$
@@
$M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}}(\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,v_{1},\cdots,v_{i}))\cong M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}}(\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,v_{1},\cdots,v_{i})).$
@@ 2.3. Hochschild homology with $\operatorname{\mathbb{F}}_{p}$-coefficients
of graded $\operatorname{\mathbb{F}}_{p}$-algebras, where $|\lambda_{i}^{\prime}|=2i+1$ by [MS93, Remark 4.3]. We write $\lambda_{j}\coloneqq\lambda_{p^{j}-1}^{\prime}$. The $\operatorname{\mathbb{E}}_{3}$-ring $\operatorname{THH}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})$ is a retract of $\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ as an $\mathbb{E}_{1}$-ring by [CM15, Corollary 1.3] and has homotopy groups
@@ 2.4. The motivic filtration on Hochschild homology
We also say a spectrum $X$ is even if $\pi_{2k-1}X=0$ for all integers $k$. We need a preliminary result, which uses the notion of even flatness due to Pstrągowski, see [PST23, Definition 4.4].
Proposition 2.13.
The $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-module $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even flat. In particular, the spectrum $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$ is even.
@@
of $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebras where
@@
By Proposition 2.13, we know that $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even. Therefore, the $p,v_{1},\cdots,v_{n}$-Bockstein spectral sequences all collapse and we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU})_{p}^{\wedge}$ is a finitely generated free $\operatorname{\mathbb{Z}}_{p}^{\wedge}[v_{1},v_{2},\cdots,v_{n}]$-module. Letting $B_{i}=\operatorname{BP}\langle i\rangle$ for $1\leq i\leq n$. We have cofiber sequences
@@
Note that the commutative diagram of $\operatorname{THH}(\operatorname{MU})$-modules
$\operatorname{MU}$$\operatorname{THH}(\operatorname{MU})$$\operatorname{BP}\langle n\rangle$$\operatorname{THH}(\operatorname{BP}\langle n\rangle)$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$}
implies that the class $v_{n+1}\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}$ acts trivially on $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})$ compatibly with the map $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})\to\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$.
Definition 2.16.
Consider the short exact sequence
$0\to M\to M\langle\varepsilon_{n+1}\rangle\to\Sigma^{2p^{n+1}-1,-1}M\to 0$
where $M=\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ induced by the cofiber sequence
$\Sigma^{2p^{n+1}-2}\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}\overset{v_{n+1}}{\longrightarrow}\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}\to\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}/v_{n+1}$
by base-change and using the observation that $v_{n+1}$ acts trivially on $M$. Then the class $\varepsilon_{n+1}$ is defined as a lift of $1\in\Sigma^{2p^{n+1}-1,-1}M$. The choice of lift $\varepsilon_{n+1}$ is unique (up to a unit) by Proposition 2.14.
@@
as well. Taking a further quotient by $v_{n+1}$ produces the result by Definition 2.16 and the discussion preceding it. ∎